Find the measure of an angle if the measure of the angle is 8 degrees less than three times the measure of its supplement.
What I have done:
x=8-3(90)
x=8-270
x=-262

Answer by algebrahouse.com(1659)

Find the measure of an angle if the measure of the angle is 8 degrees less than three times the measure of its supplement.

x = angle
180 - x = supplement

x = 3(180 - x) - 8 {measure of angle is 8 less than three times supplement}
x = 540 - 3x - 8 {used distributive property}
x = 532 - 3x {combined like terms}
4x = 532 {added 3x to both sides}
x = 133 {divided both sides by 4}
133 is the angle
www.algebrahouse.com